Fill in the boxes with the correct symbol out of >,
-8/5 [ ] -7/4

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

The LCM of the denominators 5 and 4 is 20
∴(-8/5) = [(-8 × 4)/ (5 × 4)] = (-32/20)
And (-7/4) = [(-7 × 5)/ (4 × 5)] = (-35/20)
Now,
-32 > - 35
So, (-32/20) > (- 35/20)
Hence, -8/5 [>] -7/4

​Thanks